ABCDEF is a regular hexagon. Express the vectors AC,AD,AE and AF in terms of AB and BC.

Grade:12

1 Answers

Arun
6 years ago
Dear Sumit
 
here , given AB =a

BC =b (here a and b are vectors)

EF=-b

DE=-a

and we know that AD=2BC

=2b

AD=AB+BC+CD

2b =a+b+CD

CD =b-a

and FA=-CD

=a-b

Ac=AB+BC

=a+b

AF=CD=b-a

and CF=2DE

=-2a
 
 
Regards
Arun (askIITians forum expert)